I receive my tax return of $5000 and deposit it in an interest bearing account at...

  1. I receive my tax return of $5000 and deposit it in an interest bearing account at the bank.  It receives 7% interest compounding quarterly.  What is the amount of interest that you earned after 10 years?
  1. What is the future value of $10,000,000 at 5% interest compounded quarterly over 20 years?
  1. What is the future value of $250,000 at 7.5% interest compounded semi-annually over 15 years?
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+7/(4*100))^4-1)*100
Effective Annual Rate% = 7.1859
Future value = present value*(1+ rate)^time
Future value = 5000*(1+0.071859)^10
Future value = 10007.98

interest = FV-PV = 10007.98-5000 = 5007.98

Please ask remaining parts seperately, questions are unrelated